OpenStudy (anonymous):

The difference of 32 and the quotient of a number m and 7 is 80

OpenStudy (anonymous):

32-m/7=80 m/7=32-80 m=7*48 =-336 i think its not right? am i?

OpenStudy (anonymous):

nope ots not right

OpenStudy (anonymous):

its not right you have to right an equation or an inequality

OpenStudy (anonymous):

32-m/7=80 m/7=32-80 m=7*48 =-336 i think its not right? am i?

OpenStudy (anonymous):

the eqn is m+336=0

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

32-m/7=80

OpenStudy (anonymous):

32-m/7=80